Three objects, a hoop, a disk, and a solid sphere start from the top of a hill and roll without slipping down to the bottom. Which one of the following statements is true?
A. The sphere reaches the bottom first.
B. The disk reaches the bottom first.
C. The object with the greatest mass reaches the bottom first.
D. All three reach the bottom at the same time.
E. The hoop reaches the bottom first.
1
Expert's answer
2020-10-25T18:32:24-0400

Thetime of motion

"t=\sqrt{\frac{2L}{g\sin\theta}\left(1+\frac{I}{mr^2}\right)}"

The moment of inertia

"I_{hoop}=mr^2"

"I_{disk}=\frac{1}{2}mr^2"

"I_{hoop}=\frac{2}{5}mr^2"

Hence, the time of motion is smallest for the sphere.

Answer: A. The sphere reaches the bottom first.